OpenStudy (anonymous):

find lim x->infinity h(x). f(x)=5x^3-3x^2+10x. if h(x)=f(x)/x^2. I got the answer of h(x)=5x-3+10/x. How do i find the limit?

OpenStudy (experimentx):

5x-3 = inf

OpenStudy (anonymous):

\[\lim_{x\to\infty}h(x)\mid h(x)=\frac{f(x)}{x^2}\mid f(x)=5x^3-3x^2+10x\]Let's say your solution is correct; the best thing to do now is to plug in \(\infty\) in for x. You'll just get \(\infty\) as a result.

OpenStudy (anonymous):

so when its infinity then there is no limit?